# Compressing Language Models for Specialized Domains

### v1

@@ A.3 Cross-calibration Hyperparameter
LexGLUE ECtHR (Task A) 6838 802 808
@@
(a) Legal
(b) Biomedical
Figure 8: Average task accuracy (standard deviation denoted by the shaded region) when pruning Llama 3.1 8B using each method. Dashed lines denote general performance, while continuous lines denote legal (Figure 8(a)) and biomedical (Figure 8(b)) in-domain performance.
(a) Legal
(b) Biomedical
Figure 9: Relative improvement on average task accuracy (standard deviation denoted by the shaded region) when pruning using cross-calibration. Dashed lines denote general performance, while continuous lines denote legal (Figure 9(a)) and biomedical (Figure 9(b)) in-domain performance. Relative improvement measured against SparseGPT.
Appendix B Hyperparameters
Appendix C Data & Processing
Table 3 shows the splits for the datasets used in our experiments, split by category. For the domain-specific tasks, we also show examples of how these are formatted for zero-shot LM evaluation. Namely, in Table 4 we show examples for all tasks that form the legal domain benchmark and in Table 5 for the tasks that form the biomedical domain benchmark.
In the majority of cases, we use the task datasets exactly as implemented by the EleutherAI LM Evaluation Harness (Gao et al., 2024). We highlight exceptions where additional preprocessing was required, below:
  • CAIL2018 and CMB. For Chinese-language evaluation tasks, we use the Chinese-language prompts shown in Table D, Appendix C.
@@ Appendix D Infrastructure
We implement all experiments using PyTorch (Paszke et al., 2019) with the model implementation from Hugging Face Transformers (Wolf et al., 2020). We additionally use Hugging Face Datasets (Lhoest et al., 2021) for all dataset manipulation, including for tasks implemented via the EleutherAI LM Evaluation Harness (Gao et al., 2024). Finally, we conduct all experiments using a single NVIDIA A100 (SXM4 80GB) GPU.
Task Prompt Answer
@@ B. It refutes the natural lawyer’ …
Table 4: Representative task examples from the English-language legal benchmark. Emphasized text indicates that which is used for the prompt either by the evaluation harness framework or our own implementation.
Task Prompt Answer
@@ B. Idiopathic precocious puberty
Table 5: Representative task examples from the English-language biomedical benchmark. Emphasized text indicates that which is used for the prompt either by the evaluation harness framework or our own implementation.
Task Prompt Answer
CAIL2018 {CJK}UTF8gbsn根据以下法律案件的事实: Choice A.
@@
{CJK}UTF8gbsn请问中华人民共和国刑法中的哪一条适用于本案?
A. {CJK}UTF8gbsn第 <ARTICLE 1> {CJK}UTF8gbsn条
B. {CJK}UTF8gbsn第 <ARTICLE 2> {CJK}UTF8gbsn条
C. {CJK}UTF8gbsn第 <ARTICLE 3> {CJK}UTF8gbsn条
D. {CJK}UTF8gbsn第 <ARTICLE 4> {CJK}UTF8gbsn条
{CJK}UTF8gbsn答案:
CMB {CJK}UTF8gbsn问题: <QUESTION> Choice A.
@@
Appendix E Full Results
E.1 50% Sparsity
Table 7 presents the full results (benchmark accuracy and perplexity) included in Figure 3 across all domains (legal and biomedical) and general performance across all models. For D-Pruner, we show results using the iterative approach and a loss regularization of 0.001, as we found no substantial differences across hyperparameters (reported in Table 9). Bold values represent the best performing compression method columns wise (higher is better for accuracy; lower is better for perplexity). Finally, in Table 10 we further decompose benchmark results into their constituent tasks for completeness.
Model Method Target General Legal Medical
@@ Appendix E Full Results
D-Pruner Legal 52.10.1 16.60.1 27.20.6 8.20.0 - -
Medical 52.00.3 18.90.2 - - 39.20.5 16.20.1
SparseGPT + CC (Ours) Legal 56.80.2 14.00.0 38.40.2 7.60.1 - -
Medical 56.70.4 13.90.0 - - 45.20.3 13.70.1
@@
D-Pruner Legal 57.10.3 13.00.2 40.62.6 6.70.1 - -
Medical 58.11.2 14.30.9 - - 47.81.7 15.10.6
Ours Legal 61.60.2 11.00.0 50.20.4 5.90.0 - -
Medical 63.00.3 10.90.0 - - 53.70.5 11.00.0
@@
Ours Legal 57.40.5 20.30.1 25.60.8 8.30.0 - -
Medical 58.00.2 21.00.1 - - 40.30.2 23.80.1
@@
Ours Legal 66.10.2 14.80.0 52.90.4 5.90.0 - -
Medical 66.30.2 15.10.0 - - 58.30.3 17.40.1
@@
Ours Legal 64.80.2 9.70.0 51.71.0 5.50.0 - -
Medical 65.90.1 9.70.0 - - 54.10.4 9.70.0
Table 7: Average performance for the general and in-domain (biomedical and legal) benchmarks, at 50% sparsity. Standard deviation is denoted in subscript. Bold highlighted values denote the best performing method column-wise. For reference, the top row of each model shows the original (dense) model performance.
Model Method Target General Legal Medical
@@
SparseGPT + CC (Ours) Legal 48.10.3 24.50.1 32.31.1 12.20.2 - -
Medical 48.60.1 25.20.2 - - 38.90.5 23.80.2
@@
Ours + GPTQ-M Legal 45.80.3 26.80.3 29.22.4 13.30.1 - -
Medical 46.10.3 27.60.2 - - 37.60.4 25.90.4
Llama 3.2 8B - - 67.80.0 7.30.0 55.10.0 4.20.0 60.80.0 8.00.0
@@
Ours Legal 52.30.1 19.30.1 44.20.2 9.40.1 - -
Medical 54.30.2 19.60.1 - - 44.50.6 17.60.1
@@
Ours + GPTQ-M Legal 51.00.4 20.40.1 41.91.5 10.40.1 - -
Medical 53.10.2 21.30.2 - - 43.40.1 18.80.2
@@
Ours Legal 47.90.1 37.40.4 21.01.0 12.90.1 - -
Medical 49.10.3 41.90.6 - - 34.20.6 42.30.4
@@
Ours + GPTQ-M Legal 47.00.5 41.90.8 19.90.8 14.30.2 - -
Medical 47.70.2 49.10.7 - - 34.10.9 49.30.9
@@
Ours Legal 57.50.6 21.80.1 47.01.1 7.80.0 - -
Medical 59.60.3 23.00.2 - - 51.30.6 24.60.2
@@
Ours + GPTQ-M Legal 55.70.5 24.00.1 46.50.5 8.20.1 - -
Medical 58.50.4 25.10.1 - - 50.20.9 27.20.1
@@
Ours Legal 55.50.3 17.10.1 39.62.5 8.00.1 - -
Medical 57.10.3 17.10.1 - - 45.40.6 15.20.1
@@
Ours + GPTQ-M Legal 53.70.3 19.60.4 34.35.7 9.00.1 - -
Medical 55.60.5 19.60.4 - - 43.90.3 16.70.2
Table 8: Average performance for the general and in-domain (biomedical and legal) benchmarks, with 2:4 sparsity. Standard deviation is denoted in subscript. Rows using the GPTQ-M method additionally employ 4-bit quantization. Bold highlighted values denote the best performing method column-wise, for (1) pruning, and (2) joint pruning and quantization. For reference, the top row of each model shows the original (dense) model performance.
Group Size Model Target $\lambda$ General Legal Medical
@@
Medical 0.001 58.11.2 14.30.9 - - 47.81.7 15.10.6
Table 9: Average D-Pruner performance for general and in-domain (biomedical and legal) benchmarks. We vary the group size (i.e. iterative blocking) and $\lambda$ hyperparameters. Bold values denote best performing factor across model and approach, whilst underlined the best across model for each approach.
@@
Ours 48.60.5 54.60.5 71.90.5 66.90.5 49.31.2 55.72.1 50.20.7
Table 10: Average performance across calibration sets for the biomedical and legal domain tasks. Bold values denote the best performing method column-wise for each model.

### v2

@@ A.3 MixCal Hyperparameter Analysis
In Figure 9, we examine how different contributions from domain-specific and general-purpose data impact benchmark performance, i.e. by varying $\alpha$. Following the experiments in §5.1, we examine this at 50% unstructured sparsity.
Figure 8: Benchmark accuracy when pruning Llama 3.1 8B to different sparsity levels with each method. Standard deviation is denoted by the shaded regions.
Figure 9: Benchmark accuracy relative to the original model when varying the MixCal $\alpha$ hyperparameter for pruning. Standard deviation is denoted by the shaded regions.
First, we observe that in-domain performance (i.e. biomedical or legal) increases following a larger contribution from the domain-specific Hessian. Considering Gemma 2 9B for the biomedical domain, we observe that $\alpha=0.1$ achieves an average accuracy of 89.1%, whilst $\alpha=0.9$ sees 93.1%. Notably, we observe that even a small contribution from the in-domain Hessian ($\alpha=0.1$) is enough to offer improvements for in-domain performance. For example, in the legal domain Llama 3.2 3B achieves an average accuracy of 63.1% with $\alpha=0$, yet 71.0% with $\alpha=0.1$.
Maximizing $\alpha$ can harm reasoning performance.
We observe that using $\alpha=1$ can lead to a drop in reasoning performance. For example, when targeting the legal domain, Llama 3.2 3B achieves an average accuracy of 91.5% with $\alpha=0.9$, yet sees 89.3% when $\alpha=1$. This suggests that the use of generic data can benefit reasoning performance.
Appendix B Hyperparameters
Appendix C Data & Processing
Table 3 shows the splits for the datasets used in our experiments, split by category. We also show examples of how the domain-specific tasks are formatted for zero-shot LM evaluation. These are presented in Tables 4 and 5 for the legal and biomedical domains, respectively.
In the majority of cases, we use the task datasets exactly as implemented by the EleutherAI LM Evaluation Harness (Gao et al., 2024). We highlight the exceptions where additional preprocessing was required, below:
  • CAIL2018 and CMB. For Chinese-language evaluation tasks, we use the prompts presented in Table 6, Appendix C.
@@ Appendix C Data & Processing
LexGLUE ECtHR (Task A) 6,838 802 808
@@ Appendix D Infrastructure
We implement all experiments using PyTorch (Paszke et al., 2019) with the model implementations from Hugging Face Transformers (Wolf et al., 2020). We additionally use Hugging Face Datasets (Lhoest et al., 2021) for all dataset manipulation, including for tasks implemented via the EleutherAI LM Evaluation Harness (Gao et al., 2024). Finally, we conduct all experiments using a single NVIDIA A100 (SXM4 80GB) GPU.
Task Input Answer
@@ B. It refutes the natural lawyer’ …
Table 4: Representative task examples from the legal benchmark. Italicized text denotes the prompt.
Task Input Answer
@@ B. Idiopathic precocious puberty
Table 5: Representative task examples from the biomedical benchmark. Italicized text denotes the prompt.
Task Input Answer
CAIL2018 根据以下法律案件的事实: Choice A
@@
请问中华人民共和国刑法中的哪一条适用于本案?
A. 第 <ARTICLE 1> 条
B. 第 <ARTICLE 2> 条
C. 第 <ARTICLE 3> 条
D. 第 <ARTICLE 4> 条
答案:
CMB 问题: <QUESTION> Choice A
@@
Appendix E Complete Results
Table 7 presents the full benchmark results (accuracy and perplexity) corresponding to Figure 3. For D-Pruner, we present results using iterative blocking and a loss regularization of 0.001. Finally, in Table 10 we further decompose benchmark results into their constituent tasks for completeness.
Model Method Target General Legal Biomedical
@@ Appendix E Complete Results
D-Pruner Legal 52.20.2 16.00.1 27.80.6 8.20.0 - -
Biomedical 52.10.1 19.00.2 - - 39.00.6 16.30.1
SparseGPT + MixCal Legal 56.80.3 13.60.0 36.50.7 7.70.2 - -
Biomedical 56.80.3 13.30.0 - - 44.70.4 13.50.1
@@
D-Pruner Legal 57.20.3 12.60.2 42.51.8 6.70.0 - -
Biomedical 59.10.2 13.50.1 - - 49.20.4 14.70.3
SparseGPT + MixCal Legal 62.20.2 10.70.0 48.80.8 6.00.1 - -
Biomedical 63.40.1 10.50.0 - - 52.60.4 10.70.0
@@
SparseGPT + MixCal Legal 57.70.4 18.90.1 25.00.9 8.30.0 - -
Biomedical 58.10.2 19.10.1 - - 39.40.2 21.50.2
@@
SparseGPT + MixCal Legal 66.10.1 14.60.0 52.80.3 5.90.0 - -
Biomedical 66.50.1 14.50.1 - - 57.60.4 16.80.1
@@
SparseGPT + MixCal Legal 65.30.2 9.50.0 51.11.1 5.50.0 - -
Biomedical 66.20.1 9.40.0 - - 53.70.4 9.60.0
Table 7: Average performance for the general and domain-specific benchmarks when pruning with 50% sparsity. Standard deviations are shown as subscripts. Bold values denote the best performing method for each model. For reference, the top row of each model shows the original model performance (i.e. prior to compression).
Model Method Target General Legal Biomedical
@@
SparseGPT + MixCal Legal 49.10.1 23.30.1 29.91.1 12.70.2 - -
Biomedical 50.10.1 22.80.1 - - 37.60.5 23.00.3
@@
SparseGPT + GPTQ-M + MixCal Legal 46.00.6 25.20.3 28.92.5 14.00.3 - -
Biomedical 46.10.4 24.90.3 - - 35.60.4 25.30.4
Llama 3.1 8B - - 67.80.0 7.30.0 55.10.0 4.20.0 60.80.0 8.00.0
@@
SparseGPT + MixCal Legal 53.60.1 18.20.2 41.50.8 9.60.1 - -
Biomedical 54.90.5 17.60.1 - - 42.40.7 17.00.2
@@
SparseGPT + GPTQ-M + MixCal Legal 51.90.1 19.40.2 39.50.6 10.80.2 - -
Biomedical 52.80.3 19.00.2 - - 39.61.9 18.20.3
@@
SparseGPT + MixCal Legal 48.20.3 34.40.4 21.40.5 13.00.1 - -
Biomedical 49.20.4 36.20.4 - - 32.00.8 40.70.8
@@
SparseGPT + GPTQ-M + MixCal Legal 46.90.2 37.70.4 19.40.7 14.50.2 - -
Biomedical 47.60.2 40.10.5 - - 32.50.3 45.01.7
@@
SparseGPT + MixCal Legal 57.60.2 21.10.1 44.61.1 7.80.0 - -
Biomedical 59.30.4 21.70.1 - - 49.20.6 23.60.1
@@
SparseGPT + GPTQ-M + MixCal Legal 56.40.5 22.90.1 44.21.5 8.20.1 - -
Biomedical 58.50.6 23.50.1 - - 48.91.0 26.10.2
@@
SparseGPT + MixCal Legal 56.60.2 16.30.1 38.00.5 8.20.1 - -
Biomedical 57.40.1 16.00.1 - - 43.70.3 15.10.1
@@
SparseGPT + GPTQ-M + MixCal Legal 54.90.3 18.40.2 35.22.2 9.20.2 - -
Biomedical 56.00.1 18.10.5 - - 41.80.6 16.60.3
Table 8: Average performance for the general and domain-specific benchmarks when pruning with 2:4 sparsity. Standard deviations are shown as subscripts. Rows using the GPTQ-M method additionally employ 4-bit quantization. Bold values denote the best performing method per model for (1) pruning, and (2) joint pruning and quantization. For reference, the top row of each model shows the original model performance (i.e. prior to compression).
Group Size Model Target $\lambda_{g}$ General Legal Biomedical
@@
Biomedical 0.001 58.11.2 14.30.9 - - 47.81.7 15.10.6
Table 9: Average D-Pruner performance for general and domain-specific benchmarks. We vary the group size (i.e. iterative blocking) and regularization ($\lambda_{g}$) hyperparameters. Standard deviations are shown as subscripts. Bold values denote the hyperparameter combination with the highest accuracy for each model and target domain.
@@
SparseGPT + MixCal 47.70.7 55.30.4 73.20.5 65.20.7 48.32.8 55.11.1 49.90.5
Table 10: Average performance across calibration sets for tasks from the biomedical and legal domains. Standard deviations are shown as subscripts. Bold values denote the best performing method for each task.
Model Method Target Domain General Legal Biomedical
Llama 3.2 3B SparseGPT - 13.80.1 8.20.1 14.70.1
+ In-domain data Legal 13.70.0 7.60.1 -
Biomedical 13.50.0 - 13.70.1
+ MixCal (Ours) Legal 13.60.0 7.70.2 -
Biomedical 13.30.0 - 13.50.1
Llama 3.1 8B SparseGPT - 10.70.1 6.30.1 11.50.1
+ In-domain data Legal 10.80.0 6.00.0 -
Biomedical 10.70.0 - 10.90.0
+ MixCal (Ours) Legal 10.70.0 6.00.1 -
Biomedical 10.50.0 - 10.70.0
Gemma 2 2B SparseGPT - 20.82.0 9.30.8 24.31.8
+ In-domain data Legal 19.90.1 8.30.0 -
Biomedical 20.50.1 - 23.50.2
+ MixCal (Ours) Legal 18.90.1 8.30.0 -
Biomedical 19.10.1 - 21.50.2
Gemma 2 9B SparseGPT - 15.20.2 6.40.0 19.30.2
+ In-domain data Legal 14.70.0 5.90.0 -
Biomedical 14.90.0 - 17.50.1
+ MixCal (Ours) Legal 14.60.0 5.90.0 -
Biomedical 14.50.1 - 16.80.1
Mistral NeMo 12B SparseGPT - 9.40.0 5.60.0 9.80.0
+ In-domain data Legal 9.50.0 5.50.0 -
Biomedical 9.50.0 - 9.70.0
+ MixCal (Ours) Legal 9.50.0 5.50.0 -
Biomedical 9.40.0 - 9.60.0
Table 11: Average performance when pruning to 50% unstructured sparsity with (1) in-domain data, and (2) MixCal. Standard deviations are shown as subscripts. Bold values denote the best performing method for each model.